# Lipid membrane domains control actin network viscoelasticity

### v1

@@ Results
Actin network evolves slowly on a single-phase, Ld-only membrane, with a relaxation time $\tau_{\text{Ld}}\approx 4100$ s, where $\tau$ is defined as the half life of the intermediate scattering function: $F(k,\tau)=0.5$. However, the actin network relaxes more quickly on a two-phase Lo/Ld membrane of similar overall actin density, with $\tau_{\text{Lo+Ld}}\approx 1600$ s (Fig. 3B, supplemental Video S1). Actin responds elastically to deformations, but is free to diffuse along the 2D fluid membrane, giving the composite membranes with adsorbed actin viscoelastic properties. The $\approx$2.6$\times$ acceleration in actin structural relaxation upon adding Lo inclusions to the membrane demonstrates that the domains control the viscoelastic behavior of the material.
@@
This is significant because the relaxation time is a readout of the elastic to viscous crossover of a material. Thus, in reducing the relaxation time, the lipid inclusions accelerate the onset of the viscous behavior of the actin network, making the network effectively softer over these intermediate time scales. This is a key difference with prior work on large 3D droplets immersed in a small-molecule polymer network Style et al. (2015), where the droplets stiffened the overall composite material.
@@
The images in Fig. 3A offer a qualitative explanation of the changing relaxation dynamics: the lipid domains coarsen amongst the actin network, providing a driving force for relaxation. At early times, we observe coarsening primarily via Ostwald ripening: lipids from small Lo domains dissolve in the Ld phase, and then re-condense in larger Lo domains (supplemental Video S2). The large domains exclude actin as they grow, while the small domains free up space for actin as they dissolve. At very long times, actin appears to partially unbind from the surface, giving Lo domains more freedom to coalesce (Fig. 3A, bottom, supplemental Video S2).
@@
The coupling between actin relaxation and lipid domain coarsening becomes apparent when we evolve the mean wavelength of the structure factor $a,\xi=2\pi/\langle k\rangle$ of both domains $a$ and actin $\xi$ (Fig. 3C). For the first $\approx$1500 s, $a$ and $\xi$ nearly overlap, with both growing increasingly quickly as the actin network relaxes and domains coarsen. This acceleration in growth is significant because passive lipid membrane domains grow according to a power law $a\sim t^{1/3}$ with sub-linear scaling Lifshitz and Slyozov (1961); Camley and Brown (2011). Elasticity reduces this scaling in 3D systems as it resists droplet deformation Meng et al. (2024). Here, the growing lipid domains are initially frustrated by actin elasticity, but at longer times, the actin relaxes in response to sustained deformation from the coarsening domains. At long times, the actin begins to unbind from the membrane, causing $\xi$ to plateau around $\approx$2000 s, while the domains continue to coarsen (Fig. 3C). The similar unbinding we see in the single-phase case at $\approx$4000 s (supplemental Video S1) confirms that viscoelastic 2D actin networks will always unbind from membranes at long times, but that the domains drive these dynamics much more rapidly.
@@
Dynamic modulation of membrane composition drives a non-monotonic shift in actin stiffness
@@
We observe that the Lo domains grow larger initially because cholesterol inserts preferentially into the Lo phase, consistent with the ternary phase diagram for DOPC, DPPC, and cholesterol developed by Veatch and Keller (Fig. 4A) Veatch and Keller (2003). However, after $\approx$15 minutes, the cholesterol concentration exceeds the binodal curve in the phase diagram, and the two phases begin to mix (Fig. 4A). We believe that the strong contacts with actin reduce lipid mobility, and are responsible for the relatively slow mixing time scale.
During the initial period of Lo domain growth (first $\approx$11 min), the area fraction of Lo phase $\phi_{\mathrm{Lo}}$ increases from approximately 20% to almost 50% (Fig. 4B). This compresses the Ld-tethered actin into a much smaller space, narrowing actin bundles and decreasing the mean wavelength $\xi$ (red curve in Fig. 4B). As actin is more tightly compressed, its increased elastic forces resist further deformation by the growing Lo domains. Thus, the domains deform sharply to fill all available space, increasing $L_{\mathrm{ex}}$ as sharp corners form (Fig. 4B). The dominance of actin elasticity over line tension is magnified by the increasing cholesterol composition, which reduces line tension Tsai and Feigenson (2019) and allows $L_{\mathrm{ex}}$ to grow beyond that of any sample in Fig. 2.
As the actin bundles are compressed by the growing domains, they stiffen and suppress thermal fluctuations of individual filaments. We measure the fluctuations of dilute, short “tracer” actin filaments, which are labeled with a different fluorescent dye so that they can be individually tracked amongst the bulk actin (supplemental Fig. S2). The tracers are also truncated with the actin-severing protein gelsolin, to reduce friction and entanglements with other filaments, and to ease tracking (supplemental Fig. S2). We measure the variance in position $\langle|\Delta\mathbf{x}|^{2}\rangle-\langle|\Delta\mathbf{x}|\rangle^{2}$ of n=212-254 tracer filaments over short, 20-second time-lapses, taken every few minutes during cholesterol addition (Fig. 4C, supplemental Video S3).
@@
These tracer mobility trends suggest that Lo domains stiffen the 2D composite material, seemingly contradicting our earlier results of network softening via coarsening lipid domains. However, there are key differences between changing domain composition and allowing domains of constant composition to coarsen. In a coarsening membrane, the total Ld area available to actin is constant, with growing Lo domains excluding actin as fast as shrinking domains accommodate it. Thus actin is effectively mixed by the mass-conserved transport of lipids and cholesterol.
@@
Conclusion
We present a 2D composite material with lipid domains embedded in a viscoelastic actin matrix, and use these liquid inclusions the tune the viscoelasticity of the material. Unlike prior work on 3D composite gels Style et al. (2015, 2018); Rosowski et al. (2020a, b), the characteristic sizes of the actin network and lipid domains are similar, resulting in anisotropic stresses that deform lipid domains away from a circular shape. The 2D nature of our material enhances this effect, as it aligns the actin filaments and prevents domains and filaments from “escaping” into the third dimension to dissipate their stress.
@@ Conclusion
Mammalian cells achieve exceptional plasticity when crawling through complex environments Phillips et al. (2009), all while constantly deforming, stretching, and organizing their complex plasma membrane. Reconstituted solid/liquid composite membrane materials recapture some of this complexity, and offer the possibility to develop novel interfacial materials with spatiotemporally-tunable viscoelastic properties Arnold and Takatori (2023).
Methods
Buffers
[…]
Actin and heavy meromyosin preparation
Rabbit skeletal muscle actin is purified from muscle acetone powder (Pel-Freez, catalog no: 41995-2, Lot 16743) using standard methods Spudich and Watt (1971); MacLean-Fletcher and Pollard (1980). No rabbits or other animals are directly involved in this study. Actin is stored as depolymerized globular actin (G-actin) at -80°C in G-buffer with 6% sucrose until use.
Heavy meromyosin (HMM) is purified from rabbit skeletal muscle (Pel-Freez Biologicals, Rogers, Arkansas) using standard methods Margossian and Lowey (1982). HMM is frozen in 10 mM potassium phosphate buffer (pH 7.0), 100 mM potassium chloride, 0.3 mM EGTA, 1 mM DTT, and 6% sucrose at -80°C until use.
Giant unilamellar vesicle (GUV) preparation
Giant unilamellar vesicles (GUVs) are prepared using the established electroformation method Angelova and Dimitrov (1986). Briefly, lipids are mixed with the following composition: 45.7% 1,2-dioleoyl-sn-glycero-3-phosphocholine (DOPC, Avanti catalog no: 850375P), 35% 1,2-dipalmitoyl-sn-glycero-3-phosphocholine (DPPC, Avanti catalog no: 850355C), 15% cholesterol (TCI Chemical, catalog no: C3624), 4% 1,2-dioleoyl-3-trimethylammonium-propane (DOTAP, Avanti catalog no: 890890P), and 0.3% 1,2-distearoyl-sn-glycero-3-phosphoethanolamine-N-[poly(ethylene glycol)2000-N’-carboxyfluorescein] (DSPE-PEG2k-FITC, Avanti catalog no: 810120C). This mixture of lipids and cholesterol phase-separates into liquid-ordered (Lo) domains dispersed within a liquid-disordered (Ld) continuous phase. Lipids are spread on an indium tin oxide (ITO)-coated microscope slide (Diamond Coatings, 8-12 Ohm slide) and dried under vacuum for 30 minutes.
Surface preparation
This mixture of HMM and PLL-g-PEG on the coverslip surface helps to cushion the planar lipid bilayer and create space between the membrane and glass. Lipid membrane domains experience frictional interactions with nanoscale roughness when directly supported on uncushioned glass coverslips. Thus, without the cushion of bulky HMM and PEG, lipid domains would experience kinetic arrest and fail to grow beyond the diffraction limit Honigmann et al. (2013).
Adsorbing planar lipid bilayer to treated coverslip surface
Assembling actomyosin cortex on a lipid bilayer
Planar lipid bilayers adsorbed to cover slips are incubated in 1 µM F-actin for 10 minutes at 37 °C. F-actin adsorbs uniformly to the bilayer via electrostatic attraction to DOTAP Schroer et al. (2020); Heath et al. (2013). The membrane is then cooled to room temperature, causing it to phase-separate into liquid-ordered and liquid-disordered phases. The actin then sequesters into the Ld phase. Unbound actin is washed from the cover slip with A-buffer.
Inserting cholesterol into phase-separated membrane sample
[…]
Preparing and tracking tracer actin
Every few minutes, we capture a short series of images (100 total images, 0.2 s between images) of the tracer actin filaments We track the movement of the tracer filaments using the Trackmate plugin in ImageJ Ershov et al. (2022), and calculate the variance in tracking positions over the 20-second time lapse.
Microscope for all imaging experiments
All imaging is carried out on an inverted Nikon Ti2-Eclipse microscope (Nikon Instruments) using an oil-immersion objective (Apo 100x, NA 1.45, oil). Lumencor SpectraX Multi-Line LED Light Source is used for excitation (Lumencor, Inc). Fluorescent light is spectrally filtered with emission filters (432/36, 515/30, 595/31, and 680/42; Semrock, IDEX Health and Science) and imaged on a Photometrics Prime 95 CMOS Camera (Teledyne Photometrics). Microscope images are collected using MicroManager 1.4 software Edelstein et al. (2014).
Image analysis
Actin density
Domain perimeter
Domains area binarized using adaptive thresholding. In many cases, binary thresholding merges tightly-spaced domains, which will greatly increase the excess perimeter if not corrected. To separate merged domains, we take each binary domain of more than ten pixels in size and check whether it can be easily separated into multiple distinct objects. First, the binarization threshold is increased until only 30% of the original domain pixels remain. The threshold is then iteratively lowered, 2% at a time, and the number of distinct objects greater than 15 pixels in size in the resulting binary image counted. If, after any threshold decrease, the number of binary objects in the image decreases, then the layer of pixels connecting the objects is deleted. This process continues until the original threshold is recovered, effectively severing a single row of pixels between the domains. The perimeter and area of each domain is then measured using the MATLAB Image Processing Toolbox. Finally, we account for any remaining merged domains by removing any significant outliers in $L_{\mathrm{ex}}$, as determined by a Grubbs test.
Domain curvature
Relaxation time
The relaxation time of the actin network is found using principles of differential dynamic microscopy (DDM), as described previously Cerbino and Trappe (2008); Giavazzi et al. (2009). The image structure function $D(\mathbf{k},t)$ is calculated by taking the fast Fourier transform of differences in image intensities $I$, separated by lag time $t$:
$D(k,t)=FFT\left[|I(\mathbf{x},t)-I(\mathbf{x},0)|^{2}\right].$ (4)
[…]
$F(k,t)=1-\frac{D(k,t)-B(k)}{A(k)}$ (5)
where $A(k)$ and $B(k)$ are time-independent constants. We use the known asymptotic behavior $F(k,t\rightarrow\infty)=0$ and $F(k,t=0)=1$ to find $A(k)$ and $B(k)$, using the relations $A(k)+B(k)=D(k,t\rightarrow\infty)$ and $B(k)=D(k,t=0)$. The relaxation time $\tau$ is the half life such that $F(k,\tau)=0.5$.
In Fig. 3B, we take $k=\langle k\rangle$ to be the mean wave vector of each sample, found using Eq. 3.
Domain and actin size
We find the size of lipid domains and actin using the mean wave vector of the static structure factor $S(k)$. To reduce artifacts due to photo bleaching, we first apply histogram-matched bleach correction to the time-lapse in ImageJ Miura (2020). We take the 2D fast Fourier transform (FFT) of each image and take the absolute value of the Fourier coefficients to be $S(n)=|FFT(n)|$ where $n=0,1,2,...$ is the Fourier mode. We radially bin $S(n)$ to the nearest integer mode. For example, the first diagonal point $S(n=\sqrt{2})$ would be rounded down to $n=1$ and averaged with all other coefficients $S(n=1)$. Fourier modes are converted to wave vectors using the transformation $k=2*\pi*n/L$ where $L$ is the image width. Only square images are analyzed, so that $L_{x}=L_{y}$, and thus $k_{x}=k_{y}$ for $n_{x}=n_{y}$.
Acknowledgments

### v2

@@ 3.3 Coarsening lipid domains accelerate the relaxation of viscoelastic actin networks
Actin network evolves slowly on a single-phase, Ld-only membrane, with a relaxation time $\tau_{\text{Ld}}\approx 4100$ s, where $\tau$ is defined as the half life of the intermediate scattering function: $F(k,\tau)=0.5$. However, the actin network relaxes more quickly on a two-phase Lo/Ld membrane of similar overall actin density, with $\tau_{\text{Lo+Ld}}\approx 1600$ s (Fig. 3B, supporting Video S1). While $\langle k\rangle$ is $\approx$1 µm-1 smaller in the single-phase system, the two-phase membrane relaxes faster than the single-phase across many wave vectors (supporting Fig. S3). Actin responds elastically to deformations, but is free to diffuse along the 2D fluid membrane, giving the composite membranes with adsorbed actin viscoelastic properties. The $\approx$2.6$\times$ acceleration in actin structural relaxation upon adding Lo inclusions to the membrane demonstrates that the domains control the viscoelastic behavior of the material.
@@
This is significant because the relaxation time is a readout of the elastic to viscous crossover of a material. Thus, in reducing the relaxation time, the lipid inclusions accelerate the onset of the viscous behavior of the actin network, making the network effectively softer over these intermediate time scales. This is a key difference with prior work on large 3D droplets immersed in a small-molecule polymer network 38, where the droplets stiffened the overall composite material.
@@
The images in Fig. 3A offer a qualitative explanation of the changing relaxation dynamics: the lipid domains coarsen amongst the actin network, providing a driving force for relaxation. At early times, we observe coarsening primarily via Ostwald ripening: lipids from small Lo domains dissolve in the Ld phase, and then re-condense in larger Lo domains (supporting Video S2). The large domains exclude actin as they grow, while the small domains free up space for actin as they dissolve. At very long times, actin appears to partially unbind from the surface, giving Lo domains more freedom to coalesce (Fig. 3A, bottom, supporting Video S2).
@@
The coupling between actin relaxation and lipid domain coarsening becomes apparent when we evolve the mean wavelength of the structure factor $a,\xi=2\pi/\langle k\rangle$ of both domains $a$ and actin $\xi$ (Fig. 3C). For the first $\approx$1500 s, $a$ and $\xi$ nearly overlap, with both growing increasingly quickly as the actin network relaxes and domains coarsen. This acceleration in growth is significant because passive lipid membrane domains grow according to a power law $a\sim t^{1/3}$ with sub-linear scaling 39, 40. Elasticity reduces this scaling in 3D systems as it resists droplet deformation 19.
Here, the growing lipid domains are initially frustrated by actin elasticity, but at longer times, the actin relaxes in response to sustained deformation from the coarsening domains. Similar viscoelastic relaxation has been observed in 3D fibrillar networks, in which coarsening droplets drive sporadic and abrupt plastic deformations in the network architecture 22. At long times, the actin begins to unbind from the membrane, causing $\xi$ to plateau around $\approx$2000 s, while the domains continue to coarsen (Fig. 3C). The similar unbinding we see in the single-phase case at $\approx$4000 s (supporting Video S1) confirms that viscoelastic 2D actin networks will always unbind from membranes at long times, but that the domains drive these dynamics much more rapidly.
@@
3.4 Dynamic modulation of membrane composition drives a non-monotonic shift in actin stiffness
@@
We observe that the Lo domains grow larger initially because cholesterol inserts preferentially into the Lo phase, consistent with the ternary phase diagram for DOPC, DPPC, and cholesterol developed by Veatch and Keller (Fig. 4A) 12. However, after $\approx$15 minutes, the cholesterol concentration exceeds the binodal curve in the phase diagram, and the two phases begin to mix (Fig. 4A). We believe that the strong contacts with actin reduce lipid mobility, and are responsible for the relatively slow mixing time scale.
During the initial period of Lo domain growth (first $\approx$11 min), the area fraction of Lo phase $\phi_{\mathrm{Lo}}$ increases from approximately 20% to almost 50% (Fig. 4B). This compresses the Ld-tethered actin into a much smaller space, narrowing actin bundles and decreasing the mean wavelength $\xi$ (red curve in Fig. 4B). As actin is more tightly compressed, its increased elastic forces resist further deformation by the growing Lo domains. Thus, the domains deform sharply to fill all available space, increasing $L_{\mathrm{ex}}$ as sharp corners form (Fig. 4B). The dominance of actin elasticity over line tension is magnified by the increasing cholesterol composition, which reduces line tension 45 and allows $L_{\mathrm{ex}}$ to grow beyond that of any sample in Fig. 2.
As the actin bundles are compressed by the growing domains, they stiffen and suppress thermal fluctuations of individual filaments. We measure the fluctuations of dilute, short “tracer” actin filaments, which are labeled with a different fluorescent dye so that they can be individually tracked amongst the bulk actin (supporting Fig. S4). The tracers are also truncated with the actin-severing protein gelsolin, to reduce friction and entanglements with other filaments, and to ease tracking (supporting Fig. S4). We measure the variance in position $\langle|\Delta\mathbf{x}|^{2}\rangle-\langle|\Delta\mathbf{x}|\rangle^{2}$ of n=212-254 tracer filaments over short, 20-second time-lapses, taken every few minutes during cholesterol addition (Fig. 4C, supporting Video S3).
@@
These tracer mobility trends suggest that Lo domains stiffen the 2D composite material, seemingly contradicting our earlier results of network softening via coarsening lipid domains. However, there are key differences between changing the molecular composition of the membrane and allowing domains of constant composition to nucleate and coarsen. In a coarsening membrane, the area fractions of Lo and Ld are constant, and large Lo domains grow by consuming smaller domains. The growing domains exclude actin at the same rate that the shrinking or consumed domains accommodate it. Thus actin is effectively mixed by the mass-conserved transport of lipids and cholesterol.
@@
4 Conclusions
We present a 2D composite material with lipid domains embedded in a viscoelastic actin matrix, and use these liquid inclusions the tune the viscoelasticity of the material. Unlike prior work on 3D composite gels 38, 14, 15, 16, the characteristic sizes of the actin network and lipid domains are similar, resulting in anisotropic stresses that deform lipid domains away from a circular shape. The 2D nature of our material enhances this effect, as it aligns the actin filaments and prevents domains and filaments from “escaping” into the third dimension to dissipate their stress.
@@ 4 Conclusions
Mammalian cells achieve exceptional plasticity when crawling through complex environments 3, all while constantly deforming, stretching, and organizing their complex plasma membrane. Reconstituted solid/liquid composite membrane materials recapture some of this complexity, and offer the possibility to develop novel interfacial materials with spatiotemporally-tunable viscoelastic properties 8.
{suppinfo}
Videos showing actin relaxation on both single- and two-phase membranes, actin relaxation with simultaneous domain coarsening, and tracer actin fluctuation; note providing additional details on cholesterol insertion into membrane; figure showing margins of composite membrane material; and figure showing tracer actin with different fluorescent dyes.
{acknowledgement}